OpenStudy (anonymous):

Lisa took a full coffee mug and drank 1/5 of its contents. When Lisa wasn't paying attention, Shelly refilled the mug by adding water to the remaining coffee and then stirring. Lisa drank 1/4 of this mixture. When she looked away again, Shelly refilled the mug with more water and stirred, and Lisa drank 1/3 of this mixture. When Lisa looked away a third time, Shelly refilled the mug with more water. What percentage of this final mixture was water?

OpenStudy (cwrw238):

just for arguments sake suppose capacity of cup = 100 ml (tiny cup!) after first refill amount water = 100/5 = 20 mls after 1/4 taken 15 mls remain second refill - 25 mls added - total 40 mls lisa drank 1/3 so 40 - 13 1/3 = 26 2/3 mls left add 33 1/3 for last refill = 26 2/3 + 33 1/3 = 60 mls 60 %

OpenStudy (anonymous):

im stuck after 15. how do you get 40 and the 33 1/3

OpenStudy (cwrw238):

well 1/4 of cup was drunk so 1/4 of 100 mls = 25 mls was added . 15 mls of water was there already giving 15+25 = 40 mls water

OpenStudy (anonymous):

oh i see now u put -25 not +25

OpenStudy (cwrw238):

oh i see i've got a habit of using - it meant a hyphen not a minus sign

OpenStudy (cwrw238):

it can cause confusion

OpenStudy (anonymous):

oh okay well thanks anyway

OpenStudy (cwrw238):

yw